OpenStudy (anonymous):

the volume of a bacteria cell is 2.56 times 10 to the negative 18th power in liters what is the volume of 10 to the 5th bacteria cells

OpenStudy (anonymous):

\[2.56\times 10^{-18}\times 10^5=2.56\times 10^{-13}\]